A spinner has 12 equal sections. 3 are blue. Find P(Blue).
3/12 = 1/4
How many students belong to BOTH A and B?
4
A counter is drawn and put back before the next draw. Is this with or without replacement?
With replacement
A fair coin is flipped twice. Find P(Heads then Heads).
1/2 × 1/2 = 1/4
If P(win) = 0.7, find P(not win).
A fair six-sided die is rolled. Find P(even number).
3/6 = 1/2
How many students are A in total?
10
A bag has 4 red and 3 purple counters. One red is already draw without replacement. Find P(Red on the second draw).
3/6 = 1/2
A bag has 1 red and 2 blue counters. Two are drawn without replacement. Find P(Red then Blue).
1/3 x 2/2 = 1/3
A fair coin is flipped twice. Find P(at least one Head).
1 − P(TT) = 1 − 1/4 = 3/4
A die was rolled 30 times. A 5 appeared 12 times. Find the experimental probability of rolling a 5.
12/30 = 2/5
A class has 20 students. 12 play soccer, 9 play basketball, and 4 play both. How many play neither?
3
A bag has 3 Golden and 4 Silver counters. Two are drawn WITH replacement. Find P(Gold then Gold).
3/7 x 3/7 = 9/49
A bag has 2 red and 3 blue counters. Two are drawn without replacement. Find P(one red and one blue, in any order).
(2/5 x 3/4) + (3/5 x 2/4) = 6/10 = 3/5
A bag has 5 red and 2 blue counters. Two are drawn without replacement. Find P(at least one Blue).
1 − P(RR) = 1 − (5/7 × 4/7) = 29/49